# CoLo-CAM: Class Activation Mapping for Object Co-Localization in Weakly-Labeled Unconstrained Videos

### v4

@@ Abstract
Keywords: Weakly-supervised Object Localization, Weakly-Labeled Unconstrained Videos, Co-localization, Class-activation Mapping (CAMs), Deep Learning, Convolutional Neural Networks
Figure 1: Illustration of training and inference procedures with our proposed CoLo-CAM method. The multi-frame training loss term is illustrated in the case where $n=3$ frames. Pixels are fully connected between three consecutive frames. Pixels with color green represent the foreground, while blue indicates the background. For clarity, only a few pixels are connected over a simplified white grid of pixels. Solid connections illustrate strong visual similarity, while dotted connections indicate weak similarity.
a) Training phase: Only the decoder $f$ is trained. The classifier (encoder $g$ + classification head) is pre-trained, and frozen. Our training loss combines per-frame and our multi-frame terms. Our new multi-frame loss jointly learns to localize an object across frames, assuming an object maintains similar color. The training detail is illustrated in Fig.2.
b) Inference phase: CoLo-CAM operates on a single frame and predicts a class label and bounding box localization. The notation and loss terms are described in Section 3.
@@ 1 Introduction
Techniques for automated video analysis have become important due to the recent progress of online multimedia platforms such as YouTube that provide easy access to a large number of videos [65, 70]. In this work, we focus on the task of localizing objects in videos, which plays a critical role in understanding video content and improving downstream tasks such as action recognition [40], video-based summarization [80], event detection [13], object detection [16, 30, 65], facial emotion recognition [78, 84], and visual object tracking [9, 47]. Unfortunately, the video annotations required for fully supervised localization, i.e., bounding boxes for each frame, are costly due to the large number of frames. As an alternative, weak supervision using a video class label or tag associated with the entire video is currently the most common annotation for weakly supervised video object localization (WSVOL) [33, 74]. A video class label describes the main object appearing in the video without spatiotemporal information, i.e., spatial location, and start/end time of the object’s appearance, which adds more uncertainty to the label. Additionally, unconstrained videos are often captured in the wild, with varying quality, moving objects and cameras, changing viewpoints and illumination, and decoding artifacts. Given these factors, localization from weakly-labeled unconstrained videos is a challenging task.
State-of-the-art WSVOL methods often follow a similar strategy. Localization is often formulated as an optimization problem solved over the entire video. Initial segments/proposals are generated based on visual or motion cues, which are then used to identify and refine object locations through post-processing [31, 42, 59, 70, 73, 77, 79, 81]. This process is optimized under visual appearance and motion constraints to ensure consistency. Other methods rely on co-localization/co-segmentation across videos/images via graphs that allow for a larger view and interactions between segments [15, 25, 33, 35, 74, 82]. Despite their success, these methods have several limitations. For instance, initial proposals are not necessarily discriminative, i.e., aligned with the video class, due to unsupervised region extraction using visual and motion cues (e.g., optical flow) [44, 68]. Moreover, video class labels are typically only used to cluster videos, which further limits the beneficial impact of discriminative information. Additionally, a per-video model is often optimized, leading to several issues, including scalability, inference time, and deployment. Finally, multi-stage training combines the best local solutions, which does not necessarily lead to the best global solution.
@@
To alleviate this issue while benefiting from CAM methods, the CoLo-CAM method is proposed for WSVOL (Figs. 1 and 2). This new multi-class discriminant approach leverages spatiotemporal information without constraining object movement. In particular, CoLo-CAM performs co-localization via explicit joint learning of CAMs across multiple frames in a video. Under the assumption that an object maintains a similar color locally, CoLo-CAM achieves co-localization by constraining a sequence of CAMs to be consistent by pushing them to activate similarly over pixels with a similar color. To avoid major changes in color/contrast of object appearance, and therefore maintain the color assumption, our joint learning is applied locally over consecutive frames. Unlike TCAM [5], CoLo-CAM does not assume a limited displacement of objects. Rather, objects can appear anywhere in the image, making for a more flexible method. This builds a consistent localization.
@@
Our main contributions are summarized as follows.
1) A CoLo-CAM method is introduced for WSVOL in unconstrained videos. Unlike the state-of-the-art TCAM method [5], our novel temporal training term does not assume that an object maintains a similar location in different frames. Instead, it allows the object to be located anywhere, making it more robust to motion. This is achieved by performing co-localization over a sequence of CAMs constrained to yield similar activation over pixels with a similar color. To that end, CoLo-CAM builds a fully connected graph between all pair of pixels to optimize a CRF loss using color.
2) Our extensive experiments and ablations on two challenging public datasets of unconstrained videos, YouTube-Objects [36, 59], show the merits of our proposed method and its robustness to long dependency. Most importantly, our produced CAMs become more discriminative. This translates into several advantageous properties including sharp, more complete, and less noisy activations localized more precisely around objects. Moreover, small and large objects are more accurately localized. CoLo-CAM provides a new state-of-the-art localization performance on YouTube-Objects datasets. Additionally, our results confirm the potential of discriminative learning for WSVOL tasks, in contrast to relying only on motion and visual cues.
@@ 2 Related Work
(a) Localization. Different methods exploit potential proposals as pseudo-supervision for localization [59, 81]. In [59], class-specific detectors are proposed. First, segments of coherent motion [11] are extracted, and then a spatiotemporal bounding box is fitted to each segment, forming tubes. A single tube is jointly selected by minimizing energy based on the similarity between tubes, visual homogeneity over time, and the likelihood of containing an object. The selected tubes are used as box supervision to train a per-class detector [24]. Other methods seek to discover a prominent object (foreground) in (a) video(s) [42, 38, 60]. Then, similar regions are localized and refined using visual appearance and motion consistency.
@@
(b) Segmentation. Independent spatiotemporal segments are first extracted [31, 70, 77, 79] via unsupervised methods [3, 77] or with a proposals generator [86] using pretrained detectors [86]. Then, using multiple properties such as visual appearance and motion cues, these segments are pseudo-labeled, while preserving temporal consistency. Graphs, such as the conditional random field (CRF) and GrabCut approaches [63], are often used to ensure consistency. In M-CNN [73], a fully convolutional network (FCN) is combined with motion cues to estimate pseudo-labels. Using motion segmentation, a Gaussian mixture model is exploited to estimate foreground appearance potentials [57]. Combined with the prediction of the FCN, these potentials are used to produce better segmentation pseudo-labels via a GrabCut-like method [63], and these are then used to fit the FCN. The previously mentioned segmentation methods use video class labels to cluster videos with the same class. While other methods do not rely on such supervision, the process is generally similar. Foreground regions are first estimated [19, 29, 57, 75] using motion cues [68], PCA, or VideoPCA [67]. This initial segmentation is later refined via graph methods [63].
@@
(c) Co-segmentation/Co-localization. The co-segmentation approach has been leveraged to segment similar objects across a set of videos. Typical methods rely on discerning common objects via the similarity of different features including visual appearance, motion, and shape. Using initial guessing of objects/segments, graphs, such as CRF and graph cuts, are considered to model relationships between them [15, 25, 74, 82]. In [15], the authors use relative intra-video motion extracted from dense optical flow and inter-video co-features based on a Gaussian Mixture Model (GMM). The common object, i.e., the foreground, in two videos is isolated from the background using a Markov random field (MRF). It is solved iteratively via graph cuts while accounting for a unary term based on the distance to GMMs, and a pairwise energy based on feature distance weighted by the relative motion distance between super-voxels.
Other methods aim to co-localize objects in images or videos using bounding boxes instead of segments [33, 35]. In [35], objectness [2] is used to generate proposals to search for the best candidate. Box similarity and discriminability [71] in addition to multiple object properties including visual appearance, position, and size are used as constraints to ensure consistency. Authors in [33] leverage co-saliency as before filtering out noisy bounding box proposals.
@@
(d) WSOL in still images. CAM-based methods have emerged as a dominant approach for the WSOL [17, 62] on still images. Using only global image-class labels, a DL model can be trained to classify an image and yield a discriminative heatmap per class, i.e. CAM, to localize the object in the image. Early work of these methods focused on designing variant spatial class-scores pooling layers [22, 23, 55, 87]. However, CAMs tend to under-activate by highlighting only small and the most discriminative parts of an object [17, 62], diminishing its localization performance. To overcome this, different strategies have been considered, including data augmentation on input image or deep features [7, 18, 48], as well as architectural changes [27, 43, 85]. Recently, learning via pseudo-labels has shown good results, despite its reliance on noisy labels [6, 8, 49, 50, 51, 52, 76]. Most previous methods use only forward information in a model. Other methods leverage backward information as well. These include biologically inspired methods [12, 83], and gradient [14, 26, 34, 64] or confidence/score aggregation methods [20, 53]. Recently, prompt-based methods such as CLIP-ES [45] trained using text and images allow the prediction of ROIs related to a specific input class. However, such models lack classification capability. While CAM-based methods are successful on still images, they still require adaptation to leverage the temporal information in videos. Authors in [5] considered improving the quality of pseudo-labels to train an F-CAM architecture [8]. To this end, they aggregate CAMs extracted from a sequence of frames into a single CAM that covers more ROIs. However, this led to limited success since degradation in localization performance has been reported after a sequence of 2 frames.
@@
Despite the reported success, state-of-the-art WSVOL methods still suffer different limitations. Proposal generation is a critical step often achieved through unsupervised learning using color cues and motion [31, 73, 77, 79]. Additionally, the discriminative information, i.e., video class label, is not used directly, but rather to cluster videos [33, 35, 42, 46, 81]. As such, the initial localization is vulnerable to errors, which can hinder localization performance in subsequent steps. Moreover, localization is often formulated as the solution to an optimization problem over a single/cluster of videos [33, 35, 74, 82]. This requires a model per video/cluster of videos. Therefore, the operational deployment of WSVOL methods at inference time is expensive and impractical for real-world applications, and does not scale well to large number of classes.
@@
(e) Relation to F-CAM [8] and TCAM [5]: F-CAM [8] is a WSOL method for still images. While it does not leverage temporal information, its training relies on key terms to alleviate standard CAM issues such as CAM under-activation and bloby effect [62]. It performs classification and it produces full size CAM, thanks to its U-Net like architecture [61], allowing the application of different pixel-wise losses, such as pixel pseudo-labels and CRF [72]. The recent TCAM method [5] builds on top of F-CAM. It uses the same F-CAM architecture, and its same per-frame losses. Its aim is to leverage temporal information to improve localization. This is done by improving pixel pseudo-labels via a new temporal max-pooling over low resolution CAMs from a pretrained classifier. It allows to gather relevant spatial ROIs through time. However, its localization performance declines when increasing time dependency, a consequence of the TCAM assumption about the limited movement of objects. This leads to noisy pseudo-labels and, therefore, poor localization. Our CoLo-CAM method also builds on top of F-CAM architecture, and its per-frame losses. However, the major difference is how to leverage temporal information to improve localization in videos. In this paper, a new constraint loss is proposed for joint training over a sequence of frames to perform co-localization. We explicitly constrain the output CAMs to respond similarly over objects with similar colors in a sequence of frames. This assumes objects have similar colors in adjacent frames. Meanwhile, no position assumption is made allowing objects to move freely which adds more flexibility. The interconnection created between frames/CAMs allows for knowledge transfer and correction of localization. The resulting discriminative CAMs provide a more precise object coverage and less noisy activations.
@@ 3 Proposed Approach
Let us denote by ${\mathbb{D}=\{(\bm{V},y)_{i}\}_{i=1}^{N}}$ a training set of videos, where ${\bm{V}=\{\bm{X}_{t}\}_{t=1}^{T}}$ is a video with $T$ frames, ${\bm{X}_{t}:\Omega\subset\mathbb{R}^{2}}$ is the $t$-th frame, and ${\Omega}$ is a discrete image domain. The global video class label is denoted ${y\in\{1,\cdots,K\}}$, with $K$ being the total number of classes. We assume that the label $y$ is assigned to all frames within a video. Our model follows an F-CAM model [8] which has a U-Net style architecture [61] with skip connections (Fig.1). It is composed of two parts: a standard encoder ${g}$ with parameters ${\bm{\theta^{\prime}}}$ for image classification and a standard decoder ${f}$ with parameters ${\bm{\theta}}$ for localization.

### v5

@@ Abstract
Keywords: Convolutional Neural Networks, Weakly-Supervised Video Object Localization, Unconstrained Videos, Class Activation Maps (CAMs).
@@ 1 Introduction
Techniques for automated video analysis have become important due to the recent progress of online multimedia platforms such as YouTube that provide easy access to a large number of videos. In this work, we focus on the task of localizing objects in videos, which plays a critical role in understanding video content and improving downstream tasks such as action recognition [98], video-based summarization [101], and event detection [67], object detection [66], facial emotion recognition [84, 93], and visual object tracking [45]. Unfortunately, the video annotations required for fully supervised localization, i.e., bounding boxes for each frame, are costly due to the large number of frames. As an alternative, weak supervision using a video class label or tag associated with the entire video is currently the most common annotation for weakly supervised video object localization (WSVOL) [31, 78] [^2] The WSVOL task is not to be confused with Weakly-supervised Temporal Action Localization (WTAL) task [61] as they both aim to localize an object-class in videos using only video-class label. However, WTAL task aims to predict time segment of an action class while WSVOL aims for spatial localization (bounding boxes) of objects within a frame in a video.. A video class label describes the main object appearing in the video without spatiotemporal information, i.e., spatial location, and start/end time of the object’s appearance, which adds more uncertainty to the label. Additionally, unconstrained videos are often captured in the wild, with varying quality, moving objects and cameras, changing viewpoints and illumination, and decoding artifacts. This makes localization from weakly-labeled unconstrained videos a challenging task.
State-of-the-art WSVOL methods often follow a similar strategy. Localization is often formulated as an optimization problem solved over the entire video. Initial segments/proposals are generated based on visual or motion cues, which are then used to identify and refine object locations through post-processing [38, 60, 77, 90]. This process is optimized under visual appearance and motion constraints to ensure consistency. Other methods rely on co-localization/co-segmentation across videos/images via graphs that allow for a larger view and interactions between segments [11, 31, 33, 78]. Despite their success, these methods have several limitations. For instance, initial proposals are not necessarily discriminative, i.e., aligned with the video class, due to unsupervised region extraction using visual and motion cues (e.g., optical flow). Moreover, video class labels are typically only used to cluster videos, which further limits the beneficial impact of discriminative information. Additionally, a per-video model is often optimized, leading to several issues, including scalability, inference time, and deployment. Finally, multi-stage training combines the best local solutions, which does not necessarily lead to the best global solution.
@@
To alleviate this issue while benefiting from CAM methods, the CoLo-CAM method is proposed for WSVOL (Figs. 6 and 1). This new multi-class discriminant approach leverages spatiotemporal information without constraining object movement. In particular, CoLo-CAM performs co-localization via explicit joint learning of CAMs across multiple frames in a video. Under the assumption that an object maintains a similar color locally, CoLo-CAM achieves co-localization by constraining a sequence of CAMs to be consistent by pushing them to activate similarly over pixels with a similar color. To avoid major changes in color/contrast of object appearance, and therefore maintain the color assumption, our joint learning is applied locally over consecutive frames. Unlike TCAM [3], CoLo-CAM does not assume a limited displacement of objects. Rather, objects can appear anywhere in the image, making for a more flexible method, and leading to a consistent localization.
@@
Our main contributions are summarized as follows:
(1) A CoLo-CAM method is introduced for WSVOL in unconstrained videos. Unlike the state-of-the-art TCAM method [3], our novel temporal training term does not assume that an object maintains a similar location in different frames. Instead, it allows the object to be located anywhere, making it more robust to motion. This is achieved by performing co-localization over a sequence of CAMs constrained to yield similar activation over pixels with a similar color. To that end, CoLo-CAM builds a fully connected graph between all pair of pixels to optimize a CRF loss using color.
(2) Our extensive experiments and ablations on two challenging public datasets of unconstrained videos, YouTube-Objects [60, 34], show the merits of our proposed method and its robustness to long dependency. Most importantly, our produced CAMs become more discriminative. This translates into several advantageous properties including sharp, more complete, and less noisy activations localized more precisely around objects. Moreover, small and large objects are more accurately localized. CoLo-CAM provides a new state-of-the-art localization performance on YouTube-Objects datasets. Additionally, our results confirm the potential of discriminative learning for WSVOL tasks, in contrast to relying only on motion and visual cues.
@@ 2 Related Work
Localization. Different methods exploit potential proposals as pseudo-supervision for localization [60, 90]. In [60], class-specific detectors are proposed. First, segments of coherent motion [8] are extracted, then a spatiotemporal bounding box is fitted to each segment forming tubes. A single tube is jointly selected by minimizing energy based on the similarity between tubes, visual homogeneity over time, and the likelihood to contain an object. The selected tubes are used as box supervision to train a per-class detector [20]. In [90], a DL model is trained using noisy supervision to perform segmentation and localization. Given a set of videos from the same class, bounding boxes and segmentation proposals are estimated using advanced optical flow [40]. Other methods seek to discover a prominent object (foreground) in (a) video(s) [36, 62, 38]. Then, similar regions are localized and refined using visual appearance and motion consistency. In [62], box proposals are generated in a video, and only the relevant ones are retained for building an object appearance model. Maximum a posteriori inference over an undirected chain graphical model is employed to enforce the temporal appearance consistency between adjacent frames in the same video. The method in [38] leverages two simultaneous and complementary processes for object localization: the discovery of similar objects in different videos, and the tracking of prominent regions in individual videos. Region proposals [47], in addition to appearance and motion confidence, are used to determine foreground objects [8], and to maintain temporal consistency between consecutive frames.
@@
Segmentation. Independent spatiotemporal segments are first extracted [28, 74, 83, 86] via unsupervised methods [2, 83] or with a proposals generator [97] using pretrained detectors [97]. Then, using multiple properties such as visual appearance, and motion cues, these segments are labeled, while preserving temporal consistency. Graphs, such as Conditional Random Field (CRF) and GrabCut approaches [64], are often employed to ensure consistency. In [44] deal with multi-class video segmentation is proposed, where a nearest neighbor-based label transfer between videos is exploited. First, videos are first segmented into spatiotemporal supervoxels [83], and represented by high-dimensional feature vectors using color, texture, patterns, and motion. These features are compressed using binary hashing for semantic similarity measures. A label transfer algorithm is designed using a graph that fosters label smoothness between spatiotemporal adjacent supervoxels in the same video, and between visually similar supervoxels across videos. In M-CNN [77], a fully convolutional network (FCN) is combined with motion cues to estimate pseudo-labels. Using motion segmentation, a Gaussian mixture model is exploited to estimate foreground appearance potentials [58]. Combined with the FCN prediction, these potentials are used to yield better segmentation pseudo-labels via a GrabCut-like method [64], and these are then used to fit the FCN. The previously mentioned segmentation methods use video tags to cluster videos with the same class. While other methods do not rely on such supervision, the process is generally similar. Foreground regions are first estimated [14, 27, 58, 79] using motion cues [72], or PCA, or VideoPCA [70]. This initial segmentation is later refined via graph-methods [64].
@@
Co-segmentation/Co-localization. The co-segmentation methods have been leveraged to segment similar objects across a set of videos. Typical methods rely on discerning common objects via the similarity of different features including visual appearance, motion, and shape. Using initial guessing of objects/segments, graphs, such as CRF and graph cuts, are considered to model relationships between them [11, 22, 78, 91]. In [11], the authors use relative intra-video motion extracted from dense optical flow, and inter-video co-features based on Gaussian Mixture Models (GMM). The common object, i.e., the foreground, in two videos is isolated from the background using a Markov Random Field (MRF). It is solved iteratively via graph cuts while accounting for a unary term based on the distance to GMMs, and a pairwise energy based on feature distance weighted by the relative motion distance between super-voxels. The authors in [22] pursue a similar path. The method starts by generating class-independent proposals [19]. To identify the foreground object in each frame, various object characteristics are used while accounting for inter- and intra- video coherence of the foreground. A co-selection graph is formulated as a CRF exploiting unary energy from motion (optical flow), and pairwise energy over the foreground that combines visual (histogram) and shape similarities. In [91], co-segmentation in videos is performed by sampling, tracking, and matching object proposals using graphs. It prunes away noisy segments in each video by selecting proposal tracklets that are spatially salient and temporally consistent. An iterative grouping process is used to gather objects with similar shapes and appearances (histogram) within and across videos. The final object regions obtained are used to initialize a segmentation process [24]. The authors in [78] use a pre-trained FCN to generate object-like tracklets which are linked for each object category via a graph. To define the corresponding relation between tracklets, a sub-modular optimization is formulated based on the similarities of the tracklets via object appearance, shape, and motion. To discover prominent objects in each video, tracklets are ranked based on their mutual similarities.
Other methods aim to co-localize objects in images or videos using bounding boxes instead of segments. In [33], a large number of proposed bounding boxes is generated by relying on objectness [1] with the goal being to select a single box that is more likely to contain the common object across images/videos. This is achieved by solving a quadratic problem using the Frank-Wolfe algorithm [21]. Following [75], box similarity and discriminability are used as constraints. Additionally, to ensure consistency between consecutive frames, multiple object properties including visual appearance, position, and size are used. [31] leverage co-saliency as a prior to filter out noisy bounding box proposals. The co-saliency map is derived from inter- and intra-video commonness, and total motion saliency maps. Potential proposals are used for tracklet generation, and tracklets with a high confidence score and exhibiting good spatiotemporal consistency yield the final localization.
@@
WSOL in still images. CAM-based methods have emerged as a dominant approach for the WSOL [12, 50, 63] on still images. Early works focused on designing variant spatial pooling layers [17, 18, 42, 57, 59, 71, 99, 100]. An extension to a multi-instance learning framework (MIL) has been considered [30]. However, CAMs tend to under-activate by highlighting only small and the most discriminative parts of an object [12, 63], diminishing its localization performance. To overcome this, different strategies have been considered, including data augmentation over input image or deep features [6, 13, 41, 46, 69, 82, 88, 94, 102], as well as architectural changes [25, 35, 39, 85, 87, 96]. Recently, learning via pseudo-labels shown some potential, despite its reliance on noisy labels [5, 7, 4, 48, 49, 51, 52, 81, 89, 95]. Most previous methods used only forward information in a CNN, with some models designed to leverage backward information as well. These include biologically inspired methods [9, 92], and gradient [10, 23, 32, 65] or confidence score aggregation methods [15, 54, 55, 80]. Recently, prompt-based methods such as CLIP-ES [43] trained using text and images allow the prediction of ROIs related to a specific input class. However, such models lack direct classification capability. While CAM-based methods are successful on still images, they still require adaptation to leverage the temporal information in videos. Authors in [3] considered improving the quality of pseudo-labels to train an F-CAM architecture [7]. To this end, they aggregate CAMs extracted from a sequence of frames into a single CAM that covers more ROIs. However, this led to limited success since degradation in localization performance has been reported after a sequence of 2 frames.
@@
Despite the reported success, state-of-the-art WSVOL methods still suffer different limitations. Proposal generation is a critical step often achieved through unsupervised learning using color cues and motion [77]. Additionally, the discriminative information, i.e., video class label, is not used directly, but rather to cluster videos [31, 38, 90, 33]. As such, the initial localization is vulnerable to errors, which can hinder localization performance in subsequent steps. Moreover, localization is often formulated as the solution to an optimization problem over a single/cluster of videos [31, 78, 33]. This requires a model per video/cluster of videos. Therefore, the operational deployment of WSVOL methods at inference time is expensive and impractical for real-world applications, and does not scale well to large number of classes.
@@
Relation to F-CAM [7] and TCAM [3]: F-CAM [7] is a WSOL method for still images. While it does not leverage temporal information, its training relies on key terms to alleviate standard CAM issues such as CAM under-activation and bloby effect [63]. It performs classification and it produces full size CAM, thanks to its U-Net like architecture, allowing the application of different pixel-wise losses, such as pixel pseudo-labels and CRF [76]. The recent TCAM method [3] builds on top of F-CAM. It uses the same F-CAM architecture, and its same per-frame losses. It aims to leverage temporal information to improve localization. This is done by improving pixel pseudo-labels accuracy via a new temporal max-pooling over low resolution CAMs from a pretrained classifier. It allows to gather relevant spatial ROIs through time. However, its localization performance declines when increasing time dependency, a consequence of the TCAM assumption about the limited movement of objects. This leads to noisy pseudo-labels and, therefore, poor localization. Our CoLo-CAM method also builds on top of F-CAM architecture, and its per-frame losses. However, the major difference is how to leverage temporal information to improve localization in videos. In this paper, a new constraint loss is proposed for joint training over a sequence of frames to perform co-localization. We explicitly constrain the output CAMs to respond similarly over objects with similar colors in a sequence of frames. This assumes objects have similar colors in adjacent frames. Meanwhile, no position assumption is made allowing objects to move freely which adds more flexibility. The interconnection created between frames/CAMs allows for knowledge transfer and correction of localization. The resulting discriminative CAMs provide a more precise object coverage and less noisy activations.
@@ 3 Proposed Approach
Let us denote by ${\mathbb{D}=\{(\bm{V},y)_{i}\}_{i=1}^{N}}$ a training set of videos, where ${\bm{V}=\{\bm{X}_{t}\}_{t=1}^{T}}$ is a video with $T$ frames, ${\bm{X}_{t}:\Omega\subset\mathbb{R}^{2}}$ is the $t$-th frame, and ${\Omega}$ is a discrete image domain. The global video class label is denoted ${y\in\{1,\cdots,K\}}$, with $K$ being the total number of classes. We assume that the label $y$ is assigned to all frames within a video. Our model follows an F-CAM model [7] which has a U-Net style architecture with skip connections (Fig.6). It is composed of two parts: a standard encoder ${g}$ with parameters ${\bm{\theta^{\prime}}}$ for image classification and a standard decoder ${f}$ with parameters ${\bm{\theta}}$ for localization.